OpenStudy (anonymous):

use polynomial identity list 10 pythagorean triples (x^2+y^2)^2=(x^2-y^2)^2+(2xy)^2

OpenStudy (misty1212):

pick to numbers, one even one odd, make them small we can do this in our heads

OpenStudy (misty1212):

lol pick TWO small numbers like 2 and 3 say

OpenStudy (misty1212):

let me know when you pick two

OpenStudy (vishweshshrimali5):

\[\large{(x^2 + y^2)^2 = (x^2 - y^2)^2 + (2xy)^2}\] What are pythogorean triplets? Let a, b and c be a set of pythogorean triplet then: \[\large{c^2 = a^2 + b^2}\] where: c > a and b Now, what if I choose: \[\large{x^2 + y^2 = c}\] \[\large{x^2 - y^2 = a}\] \[\large{2xy = b}\]

OpenStudy (vishweshshrimali5):

They satisfy the pythogorean triplet condition. Now, only thing you need to do is to select some set of values of x and y and evaluate a, b and c which will be your required set of pythogorean triplet

OpenStudy (vishweshshrimali5):

Lets take an example: x = 2 and y = 1: \[\large{c = x^2 + y^2 = 2^2 + 1^2 = 4 + 1 = 5}\] \[\large{a = x^2 - y^2 = 2^2 - 1^2 = 4 - 1 = 3}\] \[\large{b = 2xy = 2 * 2*1 = 4}\] So, (a,b,c) = (3,4,5) is one set of pythogorean triplet

OpenStudy (vishweshshrimali5):

Can you now find another set of pythogorean triplet ? @cabette